OpenStudy (anonymous):

Find the vertex of the graph of the function. f(x) = 4x^2 + 24x + 32

OpenStudy (jhannybean):

Basically you're taking your equation \(y=4x^2+24+32\) that has the quadratic form \(y=ax^2+bx+c\) and converting it into vertex form: \(y=a(x-h)^2+k\)

OpenStudy (jhannybean):

So to start, we can immediately see that all the coefficients are divisible by 4. So simplify your function.

OpenStudy (anonymous):

y=4x2+24+32= y=ax2+bx+c ? = y=a(x−h)2+k right?

OpenStudy (anonymous):

a=4 b=24 c=32

OpenStudy (anonymous):

y=a(x−h)2+k = = y=4(x−h)2+k

OpenStudy (jhannybean):

You got the idea somewhat, but first lets simplify this :) You'll see what I mean by "fitting the equation into the form" :D

OpenStudy (anonymous):

yeah, kinda. what i don't get is how to get x&y values to then find h&k

OpenStudy (anonymous):

ohh and thank you so much for helping me @Jhannybean

OpenStudy (jhannybean):

Simplify this: \[\large y=\frac{4}{4}x^2+\frac{24}{4}x+\frac{32}{4}\]

OpenStudy (jhannybean):

@DanTheMan99

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

sorry mi wifi was acting up @Jhannybean

OpenStudy (anonymous):

y=x2+6x+8

OpenStudy (anonymous):

y=x2+6x+8 y=ax2+bx+c ? y=a(x−h)2+k

OpenStudy (anonymous):

but how will i get the other values?

OpenStudy (jhannybean):

Alright, now that we simplified it to \(y=x^2+6x+8\), we need to use the method of completing the square. Have you learned that method yet?

OpenStudy (anonymous):

not yet

OpenStudy (jhannybean):

In the method of completing the square, we will basically create another quadratic function into the quadratic function we already have, \(y=x^2+6x+8\) It's like an inception of quadratics that will help you find the vertex. :P

OpenStudy (jhannybean):

So let's work it out slowly. First set your function equal to zero. \(y=x^2+6x+8=0\) Next, move over your constants to one side of the equation, while leaving your variables on the other side side. \(y=x^2+6x=-8\) Following so far?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

do i get x and y intercepts?

OpenStudy (jhannybean):

Once we're dne you'll have your x and y intercepts and you'll see what I mean, just follow me on this one :P Trust!

OpenStudy (jhannybean):

Now we're going to create a quadratic on the left hand side of the function by completing the square of the left hand side. Essentially we're going to `finish` the quadratic on the left hand side so it resembles the form \(ax^2+bx+c\) but right now, we've only got \(ax^2+bx\) and THEREFORE need to find ourselves a new \(c\) value.

OpenStudy (jhannybean):

\[c=\left(\frac{b}{2}\right)^2 \longrightarrow c=\left(\frac{6}{2}\right)^2 = (3)^2 = 9\]

OpenStudy (anonymous):

so c = 9

OpenStudy (jhannybean):

Yep. So since we're `adding` c to the left side, we have to respectively add it to the right side as well

OpenStudy (jhannybean):

\[y=x^2+6x\color{red}{+9}=-8\color{red}{+9}\]

OpenStudy (anonymous):

y=x2+6x+9=−8+9 y=x2+6x+9=+1

OpenStudy (anonymous):

ohh ok

OpenStudy (jhannybean):

If you're good with simplifying quadratics, you'll see that the left hand side now is a PERFECT square! :o \(x^2+6x+9 \longrightarrow (x+3)^2\)

OpenStudy (jhannybean):

Rewrite it. \[y=(x+3)^2=1\]

OpenStudy (jhannybean):

Well, you would move the 1 over to fit your new equation to fit the vertex form \(y=a(x+h)^2+k\)

OpenStudy (jhannybean):

Therefore \(y=a(x+h)^2+k \longrightarrow y=(x+3)^2-1\)

OpenStudy (jhannybean):

Do you see how it works?

OpenStudy (jhannybean):

Whoop, I wrote it wrong, \(y=a(x-h)^2+k\)** Typo.

OpenStudy (anonymous):

y=a(x−h)2+k⟶y=(x+3)2−1

OpenStudy (anonymous):

⟶y=(x+3)^2−1

OpenStudy (jhannybean):

So we know a positive can be split into 2 negatives, therefore \(y=a(x-h)^2+k \longrightarrow y=(x-(-3))^2-1\)

OpenStudy (jhannybean):

So tell me, what are your x and y intercepts? :D

OpenStudy (anonymous):

y=-1

OpenStudy (anonymous):

xintercept = (x−(−3))2−1 = 1 = (x−(−3))2

OpenStudy (jhannybean):

Im not understanding that.

OpenStudy (anonymous):

1 = (x−(−3))^2

OpenStudy (jhannybean):

\[y=a(x\color{red}{-h})^2\color{blue}{+k} \longrightarrow y=(x-\color{red}{(-3)})^2\color{blue}{-1}\]

OpenStudy (jhannybean):

To help you identify them a little better. the part in red is your x-intercept, part in blue is your y intercept

OpenStudy (anonymous):

thats much better, sorry i was confused

OpenStudy (anonymous):

\[y=a(x−h)^{2}+k⟶y=(x−(−3))^{2}−1\]

OpenStudy (anonymous):

wait so what do i do now? the intercepts?

OpenStudy (jhannybean):

You're toldto find the vertex of the graph, which I helped you find in the post above :)

OpenStudy (anonymous):

thank you @Jhannybean , sorry im a little heard-headed but i just reread it and i completely understand

OpenStudy (jhannybean):

Awesome! Glad I could help :)

OpenStudy (anonymous):

@Jhannybean sorry to bother you again

OpenStudy (anonymous):

does it mean that the answer would be "b"?

OpenStudy (anonymous):

im so confused :(

OpenStudy (anonymous):

use is formula....\[-\frac{ b }{ 2a }\]

OpenStudy (anonymous):

do i plug in a=1 and b=6??

OpenStudy (anonymous):

\[y=ax^2+bx+c\]

OpenStudy (anonymous):

y=(x−(−3))^2−1 -\[\frac{ -3 }{ 1 }\]

OpenStudy (anonymous):

-(-3/2)

OpenStudy (anonymous):

right?

OpenStudy (anonymous):



OpenStudy (anonymous):

3

OpenStudy (anonymous):

a and b is coefficient

OpenStudy (anonymous):

all the numberse

OpenStudy (anonymous):

24/8

OpenStudy (anonymous):

add a negative sign

OpenStudy (anonymous):

-3

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

@saseal

OpenStudy (anonymous):

would it be the -3,-4 or -4,-3

OpenStudy (anonymous):

lowest point is obviously the y-axis

OpenStudy (anonymous):

im so lost

OpenStudy (dan815):

dantheman

OpenStudy (dan815):

lisnten to me, do u know how transformations work with functions

OpenStudy (anonymous):

GET A LIFE DAN SRS